# Diffusion Beats Autoregressive in Data-Constrained Settings

### v6

@@ 4.5 Why do Diffusion models outperform AR models in data-constrained settings?
To better understand why diffusion models are more data-efficient than autoregressive (AR) models, we conducted a series of controlled experiments aimed at isolating the core source of diffusion’s advantage.
Our hypothesis is that masked diffusion models benefit from a form of implicit data augmentation arising from their randomized masking and denoising objective. Unlike AR models, which are trained exclusively on a fixed left-to-right factorization, diffusion models are exposed to a diverse set of conditional prediction tasks due to random masking patterns, enabling them to learn to generate tokens in varying orders.
@@
To test whether the diversity in diffusion training could be replicated in AR models via explicit augmentation, we first applied standard perturbation-based techniques during AR training. Specifically, we used: (i) attention dropout — randomly dropping 25%, 50%, or 75% of attention weights; and (ii) token masking — masking a subset of input tokens by zeroing their attention weights across all layers, while retaining the standard next-token prediction objective.
As shown in Figures 7(a) and 7(b), neither approach improved validation loss. In all cases, AR models continued to overfit and remained far behind diffusion models trained for longer epochs. All AR baselines here used 140M parameters and were trained for 50 epochs; the red line in the plots marks the best diffusion model from Figure 4(a), trained for 500 epochs.
We next investigated whether diffusion’s advantage stems from exposure to diverse token orderings. To test this, we trained AR models with varying numbers of orderings: $N=1$ denotes standard left-to-right training, while $N=k$ adds k1k{-}1 random permutations of the sequence order. All the AR models in this setting had 278M parameters and were trained for 100 epochs.
As shown in Figure 8, increasing $N$ consistently lowered validation loss and delayed overfitting. At $N=16$, the 100-epoch validation loss of AR models approached that of diffusion, suggesting that diverse orderings are indeed a key driver of diffusion’s data efficiency.
Finally, this analysis suggests a natural continuum between the two paradigms: by controlling task diversity—through masking or reordering—we could design hybrid models that interpolate between compute efficiency (AR-like) and data efficiency (diffusion-like). Exploring this continuum is a promising direction for future work. Details of our permutation process are in Section 11.
@@ 5 Discussion
We hypothesize two main contributing factors. (i) Order specialization: AR models are trained with a fixed left-to-right factorization, so every gradient update reinforces the same prediction task, allowing them to specialize effectively. In contrast, diffusion models must generalize across many random token orderings, which hinders specialization. (ii) Stronger supervision per update: In AR training, every token in a training sequence serves as a supervised target, and the causal structure enables dense gradient updates, resulting in stable, low-variance learning. Diffusion models, however, compute loss only on a subset of masked tokens, making supervision sparser per sequence, even though gradients propagate through the entire input. As a result, each update carries less direct learning signal. Arriola et al. (arriola2025block, ) show that tuning the masking schedule can help reduce gradient variance and improve training compute efficiency.

### v7

@@ 3.5 Why do Diffusion models outperform AR models in data-constrained settings?
To better understand why diffusion models are more sample-efficient than autoregressive (AR) models, we conducted a series of controlled experiments aimed at isolating the core source of diffusion’s advantage.
We first applied standard perturbation-based techniques during AR training. Specifically, we used: (i) attention dropout — randomly dropping 25%, 50%, or 75% of attention weights; and (ii) token masking — masking a subset of input tokens by zeroing their attention weights across all layers, while retaining the standard next-token prediction objective.
As shown in Figures 8(a) and 8(b) in Appendix, neither approach improved validation loss. In all cases, AR models continued to overfit and remained far behind diffusion models trained for longer epochs. All AR baselines here used 140M parameters and were trained for 50 epochs; the red line in the plots marks the best diffusion model from Figure 2(b), trained for 500 epochs.
We next investigated whether diffusion’s advantage stems from exposure to diverse token orderings. To test this, we trained AR models with varying numbers of orderings: $N=1$ denotes standard left-to-right training, while $N=k$ adds k1k{-}1 random permutations of the sequence order. All permutations were fixed prior to training, and each training batch randomly samples from these orderings. All AR models in this setting used 278M parameters and were trained for 100 epochs. As shown in Figure 7, increasing $N$ consistently lowered validation loss and delayed overfitting. At $N=16$, the 100-epoch validation loss of AR models approached that of diffusion, suggesting that diverse orderings are indeed a key driver of diffusion’s sample efficiency.
Finally, this analysis suggests a natural continuum between the two paradigms: by controlling task diversity—through masking or reordering—we could design hybrid models that interpolate between compute efficiency (AR-like) and sample efficiency (diffusion-like). Exploring this continuum is a promising direction for future work. Details of our permutation process are in Section 12.
4 Conclusion
As the availability of high-quality data plateaus, improving sample efficiency becomes essential for scaling deep learning. In this work, we show that masked diffusion models consistently outperform autoregressive (AR) models in data-constrained regimes — when training involves repeated passes over a limited dataset. We establish new scaling laws for diffusion models, revealing their ability to extract value from repeated data far beyond what AR models can achieve. These results challenge the conventional belief that AR models are universally superior and highlight diffusion models as a compelling alternative when data—not compute—is the primary bottleneck. Looking ahead, efficient use of finite data may define the next frontier in scaling deep learning models. Although the studies have been performed in the context of language models, we believe these findings should apply across any kind of sequence modeling data, such as in robotics or healthcare.
5 Acknowledgement
We thank Alexander Li for his valuable insights and detailed feedback on the manuscript. We are also grateful to Zheyang Qin for his help in improving the figures. Finally, we thank Niklas Muennighoff, Simo Ryu and Colin Raffel for their helpful comments on the final draft of the paper. We thank Stella Biderman for pointing out the inconsistent unit used in the Critical Compute equation. We thank Lucas Beyer and You Jiacheng, in suggesting the experiments in Section 3.5, that helped us explain why diffusion models would be more sample-efficient. This work was supported in part by ONR MURI N00014-22-1-2773, ONR N00014-22-1-2096, and AFOSR FA9550-23-1-0747. The results and models presented in this work also used compute resources from the National AI Research Resource Pilot, with support from NVIDIA, including NVIDIA’s DGX Cloud product and the NVIDIA AI Enterprise Software Platform.
Appendix
6 Related Work
Deep Learning in Data-Constrainted Settings.
Deep learning progress has been largely driven by the scaling of both data and compute. However, recent analyses suggest we may soon face a data bottleneck that could inhibit continued advancement [42]. In language modeling, the dominant paradigm has been autoregressive (AR) models [41, 28, 4], which are typically trained for a single epoch to maximize exposure to unique tokens [11]. In light of looming data constraints, Muennighoff et al.[21] show that AR models can still benefit from data reuse: training for up to four epochs on repeated data achieves performance nearly on par with training on fresh data, suggesting an effective strategy for improving data efficiency. In contrast, computer vision has long embraced multi-epoch training along with aggressive data augmentation—such as random cropping, flipping, and color jittering—to expand effective dataset size and improve generalization[36, 43], particularly for discriminative tasks like classification and detection. Despite these practices, data efficiency in generative modeling remains underexplored, and the trade-offs between leading paradigms such as diffusion and AR models under constrained data regimes are still poorly understood.
Diffusion-Based Language Models.
[…]
7 Additional Results
@@
In Table 3, we report downstream results using 500M unique tokens. Guided by the critical compute threshold derived in Section 3.3, we scale training to 500M unique tokens and train a 2.3B-parameter diffusion model under the predicted compute budget. The model is trained for 130 epochs, after which we observe no signs of convergence and terminate due to compute limitations. For the autoregressive (AR) model, we report results from the checkpoint that achieves the best validation loss before overfitting..
Figure 8(a) and Figure 8(b) show that standard perturbation-based techniques, such as attention dropout (25–75%) and token masking, fail to improve validation loss. Across all settings, AR models continue to overfit and lag behind diffusion models trained for longer epochs. All AR baselines use 140M parameters and are trained for 50 epochs, whereas the red line denotes the best diffusion model from Figure 2(b), trained for 500 epochs.
Table 4 reports the negative log-likelihood (NLL; lower is better) on four diverse corpora: OSCAR [25], TinyStories[8], WikiText [20], and IndustryCorpus2 EN Sub [35]. These datasets span open-domain, narrative, encyclopedic, and industry-specific text.
Figure 9 shows validation loss versus training FLOPs for autoregressive (AR) and masked diffusion models under data-constrained settings. This figure extends Figure 1 in the main paper, by including empirical validation loss for the 25M unique token regimes.
Table 3: Downstream results for the best-performing (as per validation loss) autoregressive and diffusion models trained with 500M unique tokens. We also report the random baseline for reference.
Benchmarks Random Baseline AR Diffusion
ARC-Easy [6] 25.00 43.79 45.95
BoolQ [5] 50.00 51.87 55.26
COPA [30] 50.00 67.00 64.83
HellaSwag [45] 25.00 32.28 35.33
PiQA 50.00 65.71 65.61
RACE [16] 25.00 28.28 31.44
WinoGrande XL [32] 50.00 50.61 51.51
SciQ [14] 25.00 67.82 79.13
Lambada [27] 00.00 15.07 22.30
Note: All values represent accuracy (%). Best results shown in bold.
Table 4: Downstream NLL of best diffusion and AR models at 100M unique data points.
Model Type Flops OSCAR TinyStories WikiText IndustryCorpus2
Best ARM 4.32e18 3.98 2.96 4.94 / 4.96 3.58
Best MDM 1.24e20 3.83 2.93 4.50 / 4.52 3.44
Figure 9: Pareto frontier of validation loss (negative log-likelihood) versus training FLOPs for autoregressive (AR) and diffusion models under data-constrained settings, on three different unique data settings 25M, 50M and 100M.
8 Hyperparameter details
We use the following hyperparameters: batch size of 256 sequences, AdamW optimizer with $\beta_{1}{=}0.9$, $\beta_{2}{=}0.95$, $\epsilon{=}10^{-8}$, a learning rate schedule with peak 2e-4, minimum 2e-5, 1% warm-up, cosine decay, weight decay 0.1, and gradient clipping of 1.0.
9 Model Architecture
We adopt the Megatron-DeepSpeed framework as the foundation of our implementation, upon which we build our training and evaluation setup for the masked Diffusion Model. Similar to the “extended version of the architectures” proposed in [22], our model adheres to the general transformer design while introducing several architectural modifications to better align with modern LLM practices.
Specifically, we replace absolute positional embeddings with Rotary Positional Embeddings (RoPE) [37], which improve extrapolation to longer contexts and reduce parameter count. Furthermore, we adopt the SwiGLU activation function in the MLP blocks, which has been shown to outperform standard GELU or ReLU in both convergence and downstream performance [33]. To further simplify the architecture and enhance training stability, we substitute standard LayerNorm with RMSNorm and eliminate all bias terms. These design choices are consistent with [3, 40].
To preserve the original MLP capacity while aligning with hardware-friendly parameter sizes, we compute the feed-forward hidden size $h_{f}$ as:
$h_{f}=\left\lfloor\frac{8\cdot d_{\text{model}}}{3\cdot 64}\right\rfloor\cdot 64$
This rounding scheme ensures that the FFN hidden size remains divisible by 64 while closely matching the effective dimensionality used in SwiGLU layers.
We slightly modify the parameter count estimation formula from the original:
$P=12lh^{2}\left(1+\frac{13}{12h}+\frac{V+s}{12lh}\right)$
to better reflect our revised architecture. The original formula can be decomposed into: $4lh^{2}$ (attention), $8lh^{2}$ (MLP), $13lh$ (LayerNorm and biases), and $(V+s)h$ (token and positional embeddings). After applying our architectural adjustments—namely, using a SwiGLU-based MLP of dimension $h_{f}$, switching to RoPE (eliminating $sh$), and removing bias terms—we arrive at the revised formula:
$P=4lh^{2}+3lh\cdot h_{f}+6lh+Vh$
Table 5 presents all model configurations used in our experiments along with their parameter counts.
Table 5: Model Architectures
Name param (M) d_model origin_ffw_size ffw_size kv_size n_heads n_layers
7 7.0 128 512 320 32 4 3
14 13.6 224 896 576 32 7 4
20 19.5 288 1152 768 32 7 5
35 36.6 448 1792 1152 32 7 6
44 50.7 512 2048 1344 64 8 8
57 64.8 576 2304 1536 64 9 9
74 80.5 640 2560 1664 64 10 10
90 95.0 640 2560 1664 64 10 13
106 109.6 640 2560 1664 64 10 16
117 123.6 768 3072 2048 64 12 12
140 144.8 768 3072 2048 64 12 15
163 166.1 768 3072 2048 64 12 18
175 179.2 896 3584 2368 64 14 14
196 198.3 896 3584 2368 64 14 16
217 217.5 896 3584 2368 64 14 18
251 250.8 1024 4096 2688 64 16 16
278 275.7 1024 4096 2688 64 16 18
306 300.6 1024 4096 2688 64 16 20
425 416.9 1280 5120 3392 128 10 18
489 475.6 1280 5120 3392 128 10 21
509 495.9 1408 5632 3712 128 11 18
552 534.4 1280 5120 3392 128 10 24
587 566.7 1408 5632 3712 128 11 21
632 615.3 1536 6144 4096 128 12 19
664 637.6 1408 5632 3712 128 11 24
724 700.3 1536 6144 4096 128 12 22
816 785.2 1536 6144 4096 128 12 25
893 856.4 1792 7168 4736 128 14 20
1018 971.3 1792 7168 4736 128 14 23
1143 1086.3 1792 7168 4736 128 14 26
1266 1207.6 2048 8192 5440 128 16 22
1424 1353.6 2176 8704 5760 128 17 22
1429 1358.2 2048 8192 5440 128 16 25
1593 1508.9 2048 8192 5440 128 16 28
1609 1523.2 2176 8704 5760 128 17 25
1731 1644.9 2304 9216 6144 128 18 24
1794 1692.9 2176 8704 5760 128 17 28
2007 1899.8 2304 9216 6144 128 18 28
2283 2154.7 2304 9216 6144 128 18 32
2298 2165.3 2560 10240 6784 128 20 26
2639 2478.6 2560 10240 6784 128 20 30
2980 2791.9 2560 10240 6784 128 20 34
3530 3257.0 2688 10752 7168 128 21 36
3802 3561.3 2816 11264 7488 128 22 36
4084 3879.2 2944 11776 7808 128 23 36
4516 4231.9 3072 12288 8192 128 24 36
6796 6337.4 3584 14336 9536 128 28 40
9293 8640.6 4096 16384 10880 128 32 42
11452 10889.0 4352 17408 11584 128 32 47
12295 11444.2 4608 18432 12288 128 36 44
12569 12208.7 4608 18432 12288 128 32 47
13735 13560.0 4864 19456 12928 128 32 47
14940 14905.3 4992 19968 13312 128 32 49
16183 15028.3 5120 20480 13632 128 40 47
@@ 10 Discussion
We hypothesize two main contributing factors. (i) Order specialization: AR models are trained with a fixed left-to-right factorization, so every gradient update reinforces the same prediction task, allowing them to specialize effectively. In contrast, diffusion models must generalize across many random token orderings, which hinders specialization. (ii) Stronger supervision per update: In AR training, every token in a training sequence serves as a supervised target, and the causal structure enables dense gradient updates, resulting in stable, low-variance learning. Diffusion models, however, compute loss only on a subset of masked tokens, making supervision sparser per sequence, even though gradients propagate through the entire input. As a result, each update carries less direct learning signal. Arriola et al. [1] show that tuning the masking schedule can help reduce gradient variance and improve training compute efficiency.